OpenStudy (anonymous):

Find an equation of the line through the point (1,4) and perpendicular y = 3x + 1.

OpenStudy (anonymous):

\[4= \frac{-1}{3}*1+b\\ b= \frac{13}{3}*1\\ y=\frac{-x}{3}+ \frac{13}{3}\]

OpenStudy (anonymous):

slope of equation y=3x+1 is 3 since , we have equation as y=mx+c ,where m is the slope . when we say to lines are perpendicular it means .there slope will have product =-1 let the slope of this "required" line be n. den m*n=-1 , where m =3. so slope of the required equation is =-1/3 then consider the point through which it is supposed to pass

OpenStudy (anonymous):

so y-4/x-1= slope . which is n = -1/3 now simplfy this equation \[(y-4)\div (x-1) =-1/3\]